OpenStudy (anonymous):

how do you find the limit: lim ln(1+x) x->0 -------- 1 + x

OpenStudy (anonymous):

I don't see any reason you can't just plug in 0 and evaluate

OpenStudy (anonymous):

So you'll get this: \[\frac{ \ln(1+0) }{ 1+0 }=\frac{ \ln1 }{ 1 }=\ln 1=?\]

OpenStudy (anonymous):

oh, okay I gotcha. Thank you!

OpenStudy (anonymous):

You're welcome, glad I could help